The value of Kp, for water gas reaction is 1.06×105 at 298 K. Calculate the standard state Gibbs energy change of reaction at 298 K. Where value of R is 8.314 J K-1mol-1.


1
Expert's answer
2021-05-11T05:57:17-0400

Given, "K_p=1.06\times 10^5"


"T=298K, R=8.314 JK^{-1}mol^{-1}"


The standard state Gibbs energy-


"\Delta G^{o}=-RTlnk_p"

"=-8.314\times 298\times ln (1.06\times 10^5)\\

 =2477.572\times 5.0253
\\
 =62.69KJ"